# TokenSelect: Efficient Long-Context Inference and Length Extrapolation for LLMs via Dynamic Token-Level KV Cache Selection

### v2

@@ 5.1 Selection Function
The simplest selection function is to determine the criticality of the tokens through the dot product of $\mathbf{Q}$ and $\mathbf{K}_{\text{cache}}$, then select the top-$k$ critical ones as $\mathbf{K}_{\text{select}},\mathbf{V}_{\text{select}}$. The selected indices $\mathcal{I}$ are calculated as follows:
$\mathcal{I}_{\text{topk}}=\operatorname{TopK}\left(\mathbf{Q}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right).$ (8)
However, as discussed in Sec. 4, this approach is prone to inaccuracies due to disparities in norm of attention logits between heads. To maintain independence between heads, a better approach is to have each head select the top-$k$ most critical tokens, and then determine the final selection through voting among the heads:
$\mathcal{I}_{\text{head-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\mathbb{I}\left(i\in\operatorname{TopK}\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right)\right),$
where $\mathbb{I}$ is the indicator function. Unfortunately, despite better performance, this method relies on scatter_add and multiple topk operations, resulting in low efficiency on GPUs. Additionally, the 0/1 voting ignores the relative importance of tokens for each head. Therefore, we propose a head soft vote approach that offers better performance and efficiency. Specifically, we first calculate the per-head criticality, then normalize through softmax, and sum the results for all heads:
$\mathcal{I}_{\text{head-soft-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\sigma\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right).$
Algorithm 1 Selection Cache Algorithm
0:  $\mathbf{Q}$: current Query,  $k$: number of selected tokens,     $\mathbf{C}_{Q}$: Query cache, $\mathbf{C}_{\mathcal{I}}$: selection cache,      $\theta$: similarity threshold,      $\mathcal{S}$: selection function (Eq.(10)),      $\mathbf{f}$: first query flags (default True)
0:  $\mathcal{I}$: selected indices
1:  if $\mathbf{f}$ or $\text{cos}(\mathbf{Q},\mathbf{C}_{Q})<\theta$ then
2:     $\mathcal{I}\leftarrow\mathcal{S}(\mathbf{Q},k)$
3:     $\mathbf{C}_{\mathcal{I}}\leftarrow\mathcal{I}$
4:     $\mathbf{C}_{Q}\leftarrow\mathbf{Q}$
5:     $\mathbf{f}\leftarrow\text{False}$
6:  else
7:     $\mathcal{I}\leftarrow C_{\mathcal{I}}$
8:  end if
9:  return  $\mathcal{I}$
@@ 5.2 Optimizing Selection Frequency
In the Prefill Stage, $\mathbf{Q}_{\text{prefill}}\in\mathbb{R}^{n_{\text{in}}\times d}$ is input. In long-context scenarios, the number of tokens in the user’s input sequence $n_{\text{in}}$ may reach up to 1M, making it impractical to perform selection for each Query token. Considering the similarity of consecutive Queries, we use chunk-wise token selection, inputting $\frac{1}{c}\sum_{i=1}^{c}(\mathbf{Q}_{C})_{i}$ into the selection function, where $\mathbf{Q}_{C}\in\mathbb{R}^{c\times d}$ is the Query chunk and $c$ is the chunk size. This method helps maintain the compute-intensive nature of the Prefill Stage, preventing it from becoming memory bound.
@@ 5.3 Efficient Implementation
To ensure that our proposed TokenSelect is ready for real-world applications, efficient implementation is crucial. We first analyze the time breakdown of representative block-level selective sparse attention method, InfLLM Xiao et al. (2024a). From (1)(2)(3) in Fig. 5, we can observe that although selective sparse attention can significantly reduce the complexity of attention calculations, the actual computation time is still highly dependent on the implementation. The incompatibility with efficient attention implementations such as Flash Attention has resulted in methods requiring historical attention scores Zhang et al. (2024c); Oren et al. (2024); Li et al. (2024); Xiao et al. (2024a) being difficult to be applied in real-world Web applications. Through the analysis of InfLLM’s Flash Attention-compatible version, we make several discoveries. The initial motivation for estimating token criticality at the block-level is to reduce the overhead of selection function (mainly considering dot product calculation). However, we find that dot product is not the primary performance bottleneck. Instead, a significant portion of the overhead comes from indexing the KV Cache using selected indices and making them contiguous in GPU memory, which frequently occurs during the updating of KV blocks and the concatenation of selected KV Cache. The extensive I/O required for this operation further exacerbates the memory-bound in LLMs inference. Based on this, we propose that Paged Attention is a more suitable implementation for selective sparse attention. Using Paged KV Cache management (with page size=1 for TokenSelect), we can reduce the I/O volume for selection results from the scale of all selected KV Caches $O(2kd)$ to the scale of their indices $O(k)$. However, by observing (4) in Fig. 5, we find that we encounter another bottleneck under Paged KV Cache management. Since logically contiguous KV Cache is not entirely contiguous in GPU memory, it also needs to be made contiguous before performing computational operations. To address this issue, we draw inspiration from the concept of Paged Attention and implement a Paged Dot Product Kernel using Triton Tillet et al. (2019), which significantly improves the overall efficiency of TokenSelect.
@@ 6.1 Experimental Settings
To evaluate TokenSelect’s performance on long-context inference, we use three representative datasets: InfiniteBench Zhang et al. (2024a), RULER Hsieh et al. (2024), and LongBench Bai et al. (2024). Detailed descriptions and the evaluation metrics used are provided in Appendix C.
@@
To conduct a comprehensive evaluation of TokenSelect’s performance, we carry out benchmarks on three mainstream open-source LLMs - Qwen2-7B-Instruct Yang et al. (2024a), Llama-3-8B-Instruct Grattafiori et al. (2024), and Yi-1.5-6B-Chat AI et al. (2024) - comparing against the following state-of-the-art long-context inference methods: NTK-aware scaled RoPE, SelfExtend, StreamingLLM, InfLLM and MInference. Detailed descriptions of these methods are provided in Appendix B. It is worth noting that since MInference doesn’t support length extrapolation, we use an alternative evaluation method, applying it to Llama-3-8B-Instruct-262k (Llama3 after long-text post-training). Additionally, we do not include another state-of-the-art method, QUEST, as it does not support Grouped Query Attention.
@@
[…]
6.2 Performance Comparisons
InfiniteBench.
As shown in Table 1, our TokenSelect achieves significantly superior overall performance on InfiniteBench compared to all baseline methods, even though TokenSelect uses the smallest token budget (<3K). The fact that it significantly outperforms the original models demonstrates TokenSelect’s strong length extrapolation capability. We analyze that this is due to our adoption of a fine-grained KV Cache selection strategy, while considering the equal contribution of each head to selection, which ensures that we can select most critical tokens. Observing the performance of other methods, we find that RoPE interpolation methods (NTK, SelfExtend) generally perform poorly unless used on specially trained models such as Qwen2-7B-Instruct. The better performance of Qwen2-7B-Instruct on the original model can also be attributed to this. The sparse attention method StreamingLLM, based on fixed sparse patterns, can guarantee some of the model’s capabilities, but due to discarding a large amount of long-context information, it performs poorly on retrieval-related tasks (R.PK, R.Num, R.KV). The block-level selection method InfLLM can retain more long-context information compared to StreamingLLM. However, due to its sub-optimal block-level selection, it results in lower performance on most tasks compared to TokenSelect, even though we set a larger token budget for InfLLM. It is worth noting that Yi-1.5-6B does not perform normally on the R.KV task, as it is unable to correctly recite strings like the UUID.
@@ 6.2 Performance Comparisons
LongBench.
Due to space constraints, the results of LongBench are presented in the Appendix E. Although its relatively shorter text length makes it less suitable for evaluating state-of-the-art long-context inference methods, our TokenSelect still demonstrates superior overall performance compared to most baseline methods.
Methods InfiniteBench LongBench
En.Sum En.QA Code.D Math.F R.KV Avg.
Llama-3-8B-262K 20.2 12.4 22.1 26.6 14.4 33.9
    + MInference 20.5 12.9 22.3 33.1 12.8 38.4
Ours (w/ Llama-8K) 26.9 21.3 27.4 28.2 48.4 44.0
Table 3: Comparison of different methods on post-trained models on InfiniteBench and LongBench.
Comparing to methods based on post-trained model.
In Table 3, we present the performance of the post-trained model and long-context inference method Jiang et al. (2024) based on it. It shows that even compared to length extrapolation methods requiring additional training, the training-free TokenSelect still exhibits superior performance on most tasks. Although Minference can improve the performance of the original model, it fails to reverse the negative impact of long-text post-training on shorter text tasks (LongBench).
@@
6.3 Ablation Studies
In ablation studies, we primarily analyze the impact of different Selection Functions $\mathcal{S}$ on performance. To compare the performance of different Selection Functions $\mathcal{S}$ under low token budgets (i.e., token efficiency), we maintain the 2K+512 configuration. From Table 4, we can observe that our proposed head soft vote mechanism performs significantly better across all tasks. This indicates that using the head soft vote mechanism to balance each head’s contribution to token selection results can help us avoid the domination of selection by few heads with large attention logits.
@@ 6.3 Ablation Studies
Figure 7: Computation time vs. KV Cache lengths for single chunk prefill step using Qwen2-7B-Instruct. The vertical axis represents the number of attended tokens. SDPA denotes full attention by Flashinfer (chunk size: 512).
6.4 Hyper-parameter Analysis
@@
Similarity threshold of the Selection Cache $\theta$.
Fig. 6 shows that the Selection Cache hit rate increases significantly as the similarity threshold $\theta$ decreases, converging around $\theta=0.5$. This suggests potential for further acceleration of TokenSelect’s Decode Stage by reducing $\theta$. Performance sensitivity to $\theta$ varies across tasks. While most tasks exhibit slight performance degradation with decreasing $\theta$, and R.PK in InfiniteBench shows no degradation, more challenging retrieval tasks like R.KV demonstrate significant performance deterioration. This indicates higher dynamicity requirements for token selection in these tasks.
@@ 6.5 Efficiency Comparisons
Fig. 7 demonstrates the significant acceleration of attention computation achieved by TokenSelect during long-context inference. With a KV Cache length of 1M, TokenSelect can provide up to 23.84×23.84$\times$23.84 × speedup compared to FlashInfer, which is the inference kernel library we based on. This substantial improvement is attributed to our efficient kernel design.
@@
Fig. 8 compares the end-to-end latency of TokenSelect, InfLLM, and SDPA across various tasks. TokenSelect significantly accelerates long-context inference in real-world scenarios, achieving a maximum speedup of 4.70×4.70$\times$4.70 × over SDPA and 2.28×2.28$\times$2.28 × over the state-of-the-art long-context inference method while also delivering superior overall performance.
6.6 Scaling Beyond 1 Million Context Length
To further explore TokenSelect’s performance in extreme long-context scenarios, we design an extended benchmark with different text lengths following InfiniteBench. As illustrated in Fig. 9, our TokenSelect demonstrates the ability to accurately capture critical information with a small token budget in contexts up to 2M tokens, underscoring its potential in more application scenarios.

### v4

@@ 5.1 Selection Function
The simplest selection function is to determine the criticality of the tokens via the dot product of $\mathbf{Q}$ and $\mathbf{K}_{\text{cache}}$, then select the top-$k$ as $\mathbf{K}_{\text{select}},\mathbf{V}_{\text{select}}$. The selected indices $\mathcal{I}$ are calculated as:
$\mathcal{I}_{\text{topk}}=\operatorname{TopK}\left(\mathbf{Q}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right).\vskip-2.0pt$ (5)
However, as discussed in Sec. 4, this approach is prone to inaccuracies due to disparities in norm of attention logits between heads. To maintain independence between heads, a better approach is to have each head select the top-$k$ most critical tokens, and then determine the final selection through voting among the heads, where $\mathbb{I}$ is indicator function:
$\mathcal{I}_{\text{head-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\mathbb{I}\left(i\in\operatorname{TopK}\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right)\right)$ (6)
Unfortunately, despite better performance, this method relies on scatter_add and multiple topk operations, resulting in low efficiency on GPUs. Additionally, the 0/1 voting ignores the relative importance of tokens for each head. Therefore, we propose a head soft vote approach that offers better performance and efficiency. Specifically, we first calculate the per-head criticality, then normalize through softmax, and sum the results for all heads:
$\mathcal{I}_{\text{head-soft-vote}}=\operatorname{TopK}\left(\sum_{h=1}^{H}\sigma\left(\mathbf{Q}^{h}\cdot{\mathbf{K}_{\text{cache}}^{h}}^{\top}\right)\right).\vskip-2.0pt$ (7)
@@ 5.2 Optimizing Selection Frequency
In the prefill stage, $\mathbf{Q}_{\text{prefill}}\in\mathbb{R}^{n_{\text{in}}\times d}$ is inputed. In long-context scenarios, the number of tokens in the user’s input sequence $n_{\text{in}}$ may reach up to 1M, making it impractical to perform selection for each Query token. Considering the similarity of consecutive Queries, we use chunk-wise token selection, inputting $\frac{1}{c}\sum_{i=1}^{c}(\mathbf{Q}_{C})_{i}$ into the selection function, where $\mathbf{Q}_{C}\in\mathbb{R}^{c\times d}$ is the Query chunk and $c$ is the chunk size. This method helps maintain the compute-intensive nature of the prefill stage, preventing it from becoming memory bound.
@@ 5.3 Efficient Implementation
To ready TokenSelect for real‐world use, efficient implementation is crucial. We first analyze the time breakdown of representative block-level selective sparse attention method, InfLLM Xiao et al. (2024a). From (1)(2)(3) in Fig. 5, we can observe that, despite lowering theoretical complexity, actual runtime depends heavily on implementation. The incompatibility with efficient attention implementations such as Flash Attention has resulted in methods requiring historical attention scores (e.g., H2O, TOVA, SnapKV, InfLLM) impractical in real-world serving. Analysis of InfLLM’s Flash Attention–compatible version shows that, although block-level criticality estimation aims to cut selection overhead, the dot product isn’t the main bottleneck. Instead, indexing and coalescing selected KV Cache tokens in GPU memory (HBM)—during block updates and KV Cache concatenation—incurs heavy I/O, aggravating LLM inference’s memory-bound limits. Based on this, we propose that Paged Attention is a more suitable implementation for selective sparse attention. Using paged KV Cache management (with page size=1 for TokenSelect), we can reduce the I/O volume for selection results from the scale of all selected KV Caches $O(2kd)$ to the scale of their indices $O(k)$. However, (4) in Fig. 5 reveals another bottleneck under paged KV Cache management. Since logically contiguous KV Cache is not entirely contiguous in HBM, it also needs to be made contiguous before performing selection operations. To address this issue, we design a Paged Dot Product Kernel using Triton, which significantly improves the overall efficiency of TokenSelect. The formal description of this kernel is detailed in Algorithm 2.
@@ 6.1 Experimental Settings
To evaluate TokenSelect’s performance on long-context inference, we use three representative datasets: InfiniteBench Zhang et al. (2024a), RULER Hsieh et al. (2024), and LongBench Bai et al. (2024). Detailed descriptions and the evaluation metrics used are provided in Appendix G.
@@
To conduct a comprehensive evaluation of TokenSelect’s performance, we carry out benchmarks on three mainstream open-source LLMs-Qwen2-7B-Instruct Yang et al. (2024a), Llama-3-8B-Instruct Dubey et al. (2024), and Yi-1.5-6B-Chat AI et al. (2024)-comparing against the following state-of-the-art long-context inference methods: NTK-scaled RoPE, SelfExtend, StreamingLLM, InfLLM, SnapKV, InfiniGen, QUEST, RetrievalAttention and MInference. Detailed descriptions of these methods are provided in Appendix F. It is worth noting that the methods indicated in italics lack length-extrapolation capability; thus, we evaluate them using an alternative approach, applying them to Llama-3-8B-Instruct-262k (long-text post-trained Llama-3-8B-Instruct).
@@
In all experiments in this paper, we employ greedy decoding to ensure the reliability of the results. For our TokenSelect, we implement it on SGLang Zheng et al. (2024), which is a fast serving framework based on Flasherinfer flashinfer ai . We implement our method using PyTorch Paszke et al. (2019) and Triton Tillet et al. (2019). We follow the baseline approach, including 128 initial tokens and $n_{\text{local}}$ most recent tokens in the attention computation in addition to the $k$ selected tokens. For NTK and SelfExtend, we extend the model’s context length to 128K. For StreamLLM, we set $n_{\text{local}}=\text{4K}$ . For InfLLM, we set $k=\text{4K},n_{\text{local}}=\text{4K}$. For our TokenSelect, we set $k=\text{2K},n_{\text{local}}=\text{512}$ to demonstrate our token-level KV Cache selection allows us to achieve better performance with a smaller token budget. Due to the need to demonstrate the method under different $n_{\text{local}}$ and $k$, we denote the specific token budgets in the form of $k+n_{\text{local}}$ if they differ from the aforementioned settings. For InfiniteBench and LongBench, we set the threshold $\theta$ of the Selection Cache to 0.9. We use NVIDIA A100 to conduct all experiments. When inferencing sequences over 1M tokens, we additionally employee tensor parallelism, which is transparent to our TokenSelect.
@@
(a) LongBench-GovReport.
(b) InfiniteBench-En.QA.
(c) InfiniteBench-R.PK.
(d) InfiniteBench-R.KV.
6.2 Performance Comparisons
InfiniteBench.
As shown in Table 1, our TokenSelect achieves significantly superior overall performance on InfiniteBench compared to all baseline methods, even though TokenSelect uses the smallest token budget (<3K). The fact that it significantly outperforms the original models demonstrates TokenSelect’s strong length extrapolation capability. We analyze that this is due to our adoption of a fine-grained KV Cache selection strategy, while considering the equal contribution of each head to selection, which ensures that we can select most critical tokens. Observing the performance of other methods, we find that RoPE interpolation methods (NTK, SelfExtend) generally perform poorly unless used on specially trained models such as Qwen2-7B-Instruct. The sparse attention method StreamingLLM, based on fixed sparse patterns, can guarantee some of the model’s capabilities, but due to discarding a large amount of long-context information, it performs poorly on retrieval-related tasks (R.PK, R.Num, R.KV). The block-level selection method InfLLM can retain more long-context information compared to StreamingLLM. However, due to its sub-optimal block-level selection, it results in lower performance on most tasks compared to TokenSelect, even though we set a larger token budget for InfLLM. It is worth noting that Yi-1.5-6B does not perform normally on the R.KV task, as it is unable to correctly recite strings like the UUID.
@@ 6.2 Performance Comparisons
Methods En.QA En.MC Code.D R.PK R.Num R.KV
Llama-3-8B-Instruct-262k
SDPA (128K) 9.10 68.00 19.00 100.00 100.00 17.50
SDPA (262K) 12.40 67.30 22.10 100.00 100.00 14.40
StreamingLLM (2K+512) 6.00 66.00 18.50 5.00 5.00 1.00
SnapKV (2K+512) 11.80 67.00 18.00 100.00 100.00 0.50
InfLLM (2K+512) 7.00 37.00 20.50 100.00 100.00 0.50
InfiniGen (2K+512) 7.30 57.50 17.50 100.00 99.50 0.00
QUEST (2K+512) 8.20 67.00 18.00 100.00 100.00 0.00
RetrievalAttn. (2K+512) 7.50 67.00 19.00 100.00 100.00 14.00
MInference w/ static 8.60 43.20 20.60 92.40 96.30 0.20
MInference 12.90 65.90 22.30 100.00 100.00 12.80
Ours (2k+512) 9.70 68.00 19.00 100.00 100.00 20.60
Llama-3-8B-Instruct
Ours (2k+512) 21.32 45.85 27.41 100.00 97.29 48.40
Table 3: Performance comparison with methods based-on post-trained models. Baseline performance is referenced from Jiang et al. (2024) and Liu et al. (2024a).
Comparing to methods based-on post-trained models.
In Table 3, we present a performance comparison of baseline methods that do not support length extrapolation and must be applied to long-text post-trained models. Our results show that, even compared with models undergoing costly long-text post-training and the methods applied to them, the training-free TokenSelect exhibits superior performance on most tasks. These findings further demonstrate the effectiveness of TokenSelect in long-context inference and length extrapolation.
6.3 Ablation Studies
Selection functions $\mathcal{S}$.
To compare the performance of different selection functions $\mathcal{S}$ under low token budgets (i.e., token efficiency), we maintain the 2K+512 configuration. From Table 4, we can observe that our proposed head soft vote mechanism performs significantly better across all tasks. This indicates that using the head soft vote mechanism to balance each head’s contribution to token selection results can help us avoid the domination of selection by few heads with large attention logits.
@@ 6.3 Ablation Studies
Figure 7: Computation time v.s. KV Cache lengths for single chunk prefill step using Qwen2-7B-Instruct. The vertical axis represents the number of attended tokens. SDPA denotes full attention by Flashinfer (chunk size: 512).
Similarity threshold of the Selection Cache $\theta$.
Fig. 6 shows that the Selection Cache hit rate increases significantly as the similarity threshold $\theta$ decreases, converging around $\theta=0.5$. This suggests potential for further acceleration of TokenSelect’s decode stage by reducing $\theta$. Performance sensitivity to $\theta$ varies across tasks. While most tasks exhibit slight performance degradation with decreasing $\theta$, and R.PK in InfiniteBench shows no degradation, more challenging retrieval tasks like R.KV demonstrate significant performance deterioration. This indicates higher dynamicity requirements for token selection in these tasks. Owing to the limited generation lengths in current long-context inference benchmarks, we cannot yet precisely quantify the end-to-end speedup provided by the Selection Cache. Nonetheless, for a 7B-parameter model operating on 128K-token sequences, each cache hit reduces per-step latency by approximately 0.5 ms. For more detailed performance comparisons under different $\theta$, see Table 9 of Appendix I.
@@ 6.4 Efficiency Comparisons
Fig. 7 demonstrates the significant acceleration of attention computation achieved by TokenSelect during long-context inference. With a KV Cache length of 1M, TokenSelect can provide up to $23.84\times$ speedup compared to FlashInfer, which is the inference kernel library we based on. This substantial improvement is attributed to our efficient kernel design.
@@
Fig. 8 compares the end-to-end latency of TokenSelect, InfLLM, and SDPA across various tasks. TokenSelect significantly accelerates long-context inference in real-world scenarios, achieving a maximum speedup of $4.70\times$ over SDPA and $2.28\times$ over the state-of-the-art long-context inference method while also delivering superior overall performance.
7 Conclusion
In this paper, we introduces TokenSelect, a training-free approach for efficient long-context inference and length extrapolation. TokenSelect addresses the two major challenges faced by LLMs in processing long texts: the context length limitation from pre-training and the computational complexity of attention. This is achieved through a novel token-level selective sparse attention mechanism. Experimental results demonstrate that TokenSelect can achieve up to $23.84\times$ speedup in attention computation and up to $2.28\times$ acceleration in end-to-end inference latency, while exhibiting superior performance across multiple long-context benchmarks.
8 Limitations
[…]
Acknowledgment
This work was supported in part by the National Key R&D Program of China (Grant No.2023YFF0725001), in part by the National Natural Science Foundation of China (Grant No.92370204), in part by the Guangdong Basic and Applied Basic Research Foundation (Grant No.2023B1515120057), in part by the Education Bureau of Guangzhou Municipality.
Appendix A Formal Description of Algorithms
In Sec. 5.2, we propose the Selection Cache, which shares selection results among similar Queries to reduce selection frequency without sacrificing performance. Formally, it is defined as follows:
Algorithm 1 Selection Cache Algorithm
[…]
0:  $\mathcal{I}\in\{0,\dots,N-1\}^{k}$: indices of $k$ selected tokens
1:if $\mathbf{f}$ or $\text{cos}(\mathbf{Q},\mathbf{C}_{Q})<\theta$ then
2:  $\mathcal{I}\leftarrow\mathcal{S}(\mathbf{Q},k)$
3:  $\mathbf{C}_{\mathcal{I}}\leftarrow\mathcal{I}$
4:  $\mathbf{C}_{Q}\leftarrow\mathbf{Q}$
5:  $\mathbf{f}\leftarrow\text{False}$
6:else
7:  $\mathcal{I}\leftarrow C_{\mathcal{I}}$
8:end if
9:return $\mathcal{I}$
In Sec. 5.3, we propose the Paged Dot Product Kernel to efficiently perform token-level per-head criticality estimation under the paged KV‐cache management by significantly reducing I/O between HBM and SRAM. Formally, it is defined as follows:
Algorithm 2 Paged Dot Product Kernel
[…]
2:for all $h=0,\dots,H-1$ in parallel do
3:  $q\leftarrow\mathbf{Q}[h,:]$ {to SRAM}
4:  $h_{\text{kv}}\leftarrow h\bmod H_{\text{kv}}$
5:  for all $b=0,\dots,N-1$ in parallel do
6:   $t_{0}\leftarrow b\times B$
7:   $L\leftarrow\min\!\bigl(B,\;T-t_{0}\bigr)$
8:   for $j=0,\dots,L-1$ do
9:    $idx\leftarrow\mathbf{I}[\,t_{0}+j\,]$ {to SRAM}
10:    $k\leftarrow\mathbf{K}[\,idx,\,h_{\text{kv}},:\,]$ {to SRAM}
11:    $s\leftarrow\langle q,\,k\rangle$ {in SRAM}
12:    $\mathbf{S}[\,h,\,t_{0}+j\,]\leftarrow s$ {to HBM}
13:   end for
14:  end for
15:end for
16:return $\mathbf{S}$
Appendix B Scalability of TokenSelect
B.1 Scaling Beyond 1 Million Context Length
To further explore TokenSelect’s performance in extreme long-context scenarios, we design an extended benchmark with different text lengths following InfiniteBench. As illustrated in the Fig. 9, our TokenSelect demonstrates the ability to accurately capture critical information with a small token budget in contexts up to 2M tokens, underscoring its potential in more application scenarios.